# How many points converge to a heteroclinic network in an aperiodic way?

### v1

@@ 1. Introduction
To be more specific, consider a dynamical system on $\mathbb{R}^{d}$ defined by a differential equation $\dot{x}=f(x)$, $x\in\mathbb{R}^{d}$, which induces a sufficiently smooth flow $\phi_{t}$. For a hyperbolic equilibrium $\xi$ let $W^{\textrm{u}}(\xi)$, $W^{\textrm{s}}(\xi)$ denote the usual unstable/stable invariant manifolds, respectively. A cycle C in phase space of length $L$ consists of finitely many distinct hyperbolic equilibria $\xi_{1},\dotsc,\xi_{L}\in\mathbb{R}^{d}$ and connecting trajectories $\gamma_{l}\subset W^{\textrm{u}}(\xi_{l})\cap W^{\textrm{s}}(\xi_{l+1})$, where indices $\l$ are taken modulo $L$. A cycle is homoclinic if $L=1$ and heteroclinic if $L>1$. We identify a cycle C with its invariant set, that is, $\textsf{{C}}=\bigcup_{l=1}^{L}(\left\{\xi_{l}\right\}\cup\gamma_{l})$.
@@
Multiple cycles can make up larger structures in phase space: A network N in phase space is the connected finite union of two or more distinct cycles [^1] Note that we do not preclude the existence of infinitely many connecting trajectories between equilibria $\xi_{p}$ and $\xi_{p^{\prime}}$, for example, if $\dim(W^{\textrm{u}}(\xi_{p})\cap W^{\textrm{s}}(\xi_{p^{\prime}}))>1$. We rather restrict attention to a finite collection of connecting trajectories.. Its nodes are hyperbolic equilibria $\xi_{p}$, $p\in\left\{1,\dotsc,P\right\}$ connected by trajectories $\gamma_{q}$, $q\in\left\{1,\dotsc,Q\right\}$. In particular, if $\Xi=\left\{\xi_{p}\right\}$ is the set of nodes and $\alpha,\omega$ denote the usual limit sets, we have for any $x\in\gamma_{q}$ that $\alpha(x),\omega(x)\in\Xi$. The network as an invariant set, that is, $\textsf{{N}}=\bigcup_{p=1}^{P}\left\{\xi_{p}\right\}\cup\bigcup_{q=1}^{Q}\gamma_{q}$, may also be f.a.s.
@@
First, one may consider the stability of constituent cycles. For example, Postlethwaite and Dawes [14] observed “irregular cycling” dynamics close to a heteroclinic network and obtained restrictions for periodic cycling. More recently, Podvigina [15] extended the notion of and criteria for stability of cycles to omnicycles, periodic concatenations of cycles, making it possible to compute the stability properties explicitly in the same way as for cycles.
Second, there is the concept of switching near a homo-/heteroclinic network [16, 17, 18]. Roughly speaking, switching means that for any arbitrarily small neighborhood of the network and for any given way to follow the connections in the network, there is a trajectory within this neighborhood that does exactly that. One way to prove the existence of switching is to show that there is an invariant set where the dynamics are conjugated to a full shift of a suitable set of symbols [16]. In other words, in such examples any path around the network may be followed, but the trajectories doing so are not necessarily attracted to the network.
Consequently, there is a gap between stability of cycling dynamics and switching. On the one hand, numerical simulations seem to indicate that there may be trajectories that approach a network in an aperiodic way [14, 7, 15]—but it is not clear whether for any aperiodic path there are trajectories following it and, if so, how many initial conditions follow any given path. On the other hand, there are well-studied examples where there are obstructions to switching, such as the Kirk–Silber network [19] and more general networks [20]: There are parameter values for which trajectories may wind around one cycle arbitrarily often but as soon as they switch over to another cycle, they will stay there indefinitely.
In this note we discuss stability properties of arbitrary sequences of transitions between equilibria in a network N with a particular focus on aperiodic sequences. Section 2 introduces basic terminology on networks and sequences as well as the concept of switching. In Section 3, we then define a notion of fragmentary asymptotic stability for sequences of connections in a network, whether periodic or aperiodic. We show that our definition is compatible with the existing notions of fragmentary asymptotic stability for invariant sets. In Section 4 we discuss some consequences for aperiodic sequences. In particular, we show that if there is one f.a.s. aperiodic sequence, then there are countably infinitely many f.a.s. aperiodic sequences (Proposition 4.1), but there cannot be uncountably many f.a.s. aperiodic sequences (Proposition 4.3). Finally, in Section 5 we discuss concrete examples where aperiodicity does and does not play a role. We finish with a discussion and some open questions in Section 6.
@@ 2.1. Sequences on a network
Consider a network N in phase space with $P$ equilibria $\Xi=\left\{\xi_{p}\right\}$ and $Q$ heteroclinic connections $\gamma_{q}$. For any $\gamma_{q}$ and $x\in\gamma_{q}$ write $\alpha[\gamma_{q}]:=\alpha(x)$, $\omega[\gamma_{q}]:=\omega(x)$, so that $\gamma_{q}$ is a connection between the equilibria $\alpha[\gamma_{q}]$ and $\omega[\gamma_{q}]$.
@@
Sequences describe itineraries along the network N. If for each pair $(\xi_{p},\xi_{p^{\prime}})$ there is at most one connecting trajectory $\gamma_{q}$ in N with $\alpha[\gamma_{q}]=\xi_{p}$ and $\alpha[\gamma_{q}]=\xi_{p^{\prime}}$, then one may identify a sequence $\mathbf{q}=q_{1}q_{2}\ldots$ of connections with a sequence $\mathbf{p}=p_{1}p_{2}\ldots\in\left\{1,\dotsc,P\right\}^{\mathbb{N}}$ of equilibria such that $\xi_{p_{k}}=\alpha[\gamma_{q_{k}}]$, $\xi_{p_{k+1}}=\omega[\gamma_{q_{k}}]$.
@@
an invariant set containing all connections in $\mathbf{q}$ as well as the equilibria that are their endpoints. One may define $\Gamma(\mathbf{p})$ in an analogous way.
@@
A (homoclinic or heteroclinic) cycle as an invariant set C can be associated with a periodic sequence $\mathbf{q}$. In this case $\Gamma(\mathbf{q})=\textsf{{C}}$ is topologically a circle. Similarly, omnicycles are associated with a periodic sequence. Conversely, for any periodic sequence $\mathbf{q}$ the invariant set $\Gamma(\mathbf{q})$ is, in general, topologically a union of circles.
Moving along a sequence now corresponds to a shift operation.
@@
The Kirk–Silber network [19] contains four equilibria $\xi_{1},\dotsc,\xi_{4}$ and it is the union of two distinct cycles with one joint connection, cf. Figure 1. If $[\xi_{p}\to\xi_{p^{\prime}}]$ denotes the heteroclinic connection from $\xi_{p}$ to $\xi_{p^{\prime}}$, then the first cycle has connections $\gamma_{1}=[\xi_{1}\to\xi_{2}]$, $\gamma_{2}=[\xi_{2}\to\xi_{3}]$, $\gamma_{3}=[\xi_{3}\to\xi_{1}]$ and the second cycle has $\gamma_{1}=[\xi_{1}\to\xi_{2}],\gamma_{4}=[\xi_{2}\to\xi_{4}]$, $\gamma_{5}=[\xi_{4}\to\xi_{1}]$. The only relevant information is what happens at the distribution node $\xi_{2}$: All sequences $\mathbf{q}$ that start with the joint connection $\gamma_{1}$ are concatenations of the symbols $\mathrm{A}:=123$, $\mathrm{B}:=145$, i.e., they can be written compactly as sequences on the symbols $\left\{\mathrm{A},\mathrm{B}\right\}$. (For sequences of equilibria $\mathbf{p}$, the symbols are $\mathrm{A}:=123$, $\mathrm{B}:=124$.) The shift $\Sigma$ induces a shift operator $\widetilde{\Sigma}:\left\{\mathrm{A},\mathrm{B}\right\}^{\mathbb{N}}\to\left\{\mathrm{A},\mathrm{B}\right\}^{\mathbb{N}}$ that corresponds to a full transition around one of the two cycles.
[…]
@@ 2.2. Switching near a network
Sequences may be used to encode the dynamics near a network N. Let $U_{\textsf{{N}}}$ be a neighborhood of the network and let $U_{p}$ be a neighborhood of $\xi_{p}$. For each connection $\gamma_{q}$ in N consider a point $\zeta_{q}\in\gamma_{q}$ and let $V_{q}$ be a neighborhood of $\zeta_{q}$. Assume that all neighborhoods $U_{p},V_{q}$ are disjoint.
@@
(2) $\mathcal{S}_{\delta}({\mathbf{q}})$ =$=\left\{\,x\in B_{\delta}(\textsf{{N}})\,\left|\;\phi_{t}(x)\text{ follows }\mathbf{q}\text{ with }U_{\textsf{{N}}}=B_{\delta}(\textsf{{N}})\right.\right\}$
@@
It follows from the definition above that $x\in\mathcal{S}_{\delta}(\mathbf{q})$ implies $x\in\mathcal{S}_{\delta}(\Sigma\mathbf{q})$. This can be seen by applying a suitable shift to the two time series $t_{k},t^{\prime}_{k}$. Hence, $\mathcal{S}_{\delta}(\mathbf{q})\subset\mathcal{S}_{\delta}(\Sigma\mathbf{q})$. In this sense, the sequence that is realized by a given trajectory near the network is unique only up to shifting. In particular, since $t_{k},t^{\prime}_{k}\geq 0$ above, a maximal sequence realized by a given trajectory can be obtained by checking near which connection this trajectory first enters $B_{\delta}(\textsf{{N}})$ permanently, i.e., it remains in $B_{\delta}(\textsf{{N}})$ for all future times.
Equivalenty, a single trajectory realizes a countable number of sequences on the network, all of which are related by the shift operator $\Sigma$. This number is finite if the maximal sequence realized by this trajectory is preperiodic.
@@
There is (infinite) switching near the network N if in any $\delta$-neighborhood of N, every sequence $\mathbf{q}$ is realized by at least one trajectory, that is, for any $\delta>0$ and all $\mathbf{q}$ we have $\mathcal{S}_{\delta}({\mathbf{q}})\neq\emptyset$.
@@
When restricting to equilibria with real eigenvalues, it was shown recently in [20] that for a large class of networks infinite switching is impossible and even finite switching is restricted in the sense that for every network there is a bound $K\in\mathbb{N}$ such that if every finite sequence of length $k$ is realized in a network, then $k<K$.
@@ 3. Asymptotic stability of sequences
For a given sequence $\mathbf{q}$ on a network N, we are interested in how many initial conditions $x$ lead to trajectories $\phi_{t}(x)$ that not only stay close to N but rather approach N as $t\to\infty$. From now on, let $\mu$ denote the Lebesgue measure on $\mathbb{R}^{d}$.
@@ 3.1. Fragmentary asymptotic stability of sequences
(3) $\mathcal{A}_{\delta}({\mathbf{q}})$ :=$:=\big{\{}x\in\mathcal{S}_{\delta}({\mathbf{q}})\,\big{|}\;\lim_{t\to\infty}\textnormal{d}(\textsf{{N}},\phi_{t}(x))=0\big{\}}.$
@@
Typically, stability is defined for sets which are (forward) invariant with respect to the flow $\phi_{t}$. Our set $\mathcal{A}_{\delta}({\mathbf{q}})$ is not forward invariant: For a point $x\in\mathcal{A}_{\delta}({\mathbf{q}})$ and $t>0$ sufficiently large we have $\phi_{t}(x)\in\mathcal{A}_{\delta}(\Sigma^{K(t)}\mathbf{q})$ for some $K(t)>0$, but usually $\phi_{t}(x)\not\in\mathcal{A}_{\delta}(\mathbf{q})$. A concrete example can be seen in Figure 2 for $\mathbf{q}=134\dotsb$. Take a point $x$ on the blue trajectory in a $\delta$-neighborhood of $\alpha[\gamma_{1}]$, which is in $\mathcal{A}_{\delta}(134\dotsb)$. In forward time there is a $t>0$ such that $\phi_{t}(x)$ is in a $\delta$-neighborhood of $\omega[\gamma_{1}]=\alpha[\gamma_{3}]$. This means that $\phi_{t}(x)\in\mathcal{A}_{\delta}(34\dotsb)$ but in general not necessarily $\phi_{t}(x)\in\mathcal{A}_{\delta}(134\dotsb)$.
@@
(4) $\mathcal{B}_{\delta}({\mathbf{q}})$ :=$:=\left\{\,x\in B_{\delta}(\textsf{{N}})\,\left|\;\exists k\in\mathbb{N}\text{ such that }x\in\mathcal{A}_{\delta}(\Sigma^{k}\mathbf{q})\right.\right\}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{\delta}(\Sigma^{k}\mathbf{q}).$

### v2

@@ 1. Introduction
[…]
@@
Multiple cycles can make up larger structures in phase space: A network N in phase space is the connected finite union of two or more distinct cycles [^1] Note that we do not preclude the existence of infinitely many connecting trajectories between equilibria $\xi_{p}$ and $\xi_{p^{\prime}}$, for example, if $\dim(W^{\textrm{u}}(\xi_{p})\cap W^{\textrm{s}}(\xi_{p^{\prime}}))>1$. We rather restrict attention to a finite collection of connecting trajectories., possibly of different lengths. Its nodes are hyperbolic equilibria $\xi_{p}$, $p\in\left\{1,\dotsc,P\right\}$ connected by trajectories $\gamma_{q}$, $q\in\left\{1,\dotsc,Q\right\}$. In particular, if $\Xi=\left\{\xi_{p}\right\}$ is the set of nodes and $\alpha,\omega$ denote the usual limit sets, we have for any $x\in\gamma_{q}$ that $\alpha(x),\omega(x)\in\Xi$. The network as an invariant set, that is, $\textsf{{N}}=\bigcup_{p=1}^{P}\left\{\xi_{p}\right\}\cup\bigcup_{q=1}^{Q}\gamma_{q}$, may also be f.a.s.
@@
First, one may consider the stability of constituent cycles. For example, Postlethwaite and Dawes [14] observed “irregular cycling” dynamics close to a heteroclinic network and obtained restrictions for periodic cycling. More recently, Podvigina [15] extended the notion of stability of cycles to omnicycles—periodic concatenations of cycles—and gave criteria that make it possible to compute the stability properties explicitly in the same way as for cycles. In other words, omnicycles can be seen as periodic sequences of transitions along connections in a network, which yields a notion of stability for such periodic sequences (beyond any associated invariant sets).
Second, there is the concept of switching near a homo-/heteroclinic network [16, 17, 18]. Roughly speaking, switching means that for any arbitrarily small neighborhood of the network and for any given sequence to follow the connections in the network, there is a trajectory within this neighborhood that does exactly that. In some cases it is possible to prove the existence of switching by showing that there is an invariant set where the dynamics are conjugated to a full shift on a set of symbols suitably chosen to represent all possible paths around the network [16]. In such examples any sequence of network connections may be followed, but the trajectories doing so are not necessarily attracted to the network.
Consequently, there is a gap between stability of (omni)cycles and switching. On the one hand, numerical simulations indicate that there may be trajectories that approach a network in an aperiodic way [14, 7, 15]—but it is not clear whether for any aperiodic path there are trajectories following it and, if so, how many initial conditions follow any given path. On the other hand, there are well-studied examples where there are obstructions to switching, such as the Kirk–Silber network [19] and more general networks [20]: There are parameter values for which trajectories may wind around one cycle arbitrarily often but as soon as they transition to another cycle, they will stay there indefinitely.
In this note we introduce and discuss stability properties of arbitrary sequences of transitions between equilibria in a network N with a particular focus on aperiodic sequences. Section 2 introduces basic terminology on networks and sequences as well as the concept of switching. In Section 3, we then define a notion of fragmentary asymptotic stability for sequences of connections in a network, whether periodic or aperiodic. We show that our definition is compatible with the existing notions of fragmentary asymptotic stability for invariant sets. In Section 4 we discuss some consequences for aperiodic sequences. In particular, we show that if there is one f.a.s. aperiodic sequence, then there are countably infinitely many f.a.s. aperiodic sequences (Proposition 4.1), but there cannot be uncountably many f.a.s. aperiodic sequences (Proposition 4.3). Finally, in Section 5 we discuss concrete examples where aperiodicity does and does not play a role. We finish with a discussion and some open questions in Section 6.
@@ 2.1. Sequences on a network
Consider a network $\textsf{{N}}\subset\mathbb{R}^{d}$ in phase space with $P$ equilibria $\Xi=\left\{\xi_{p}\right\}$ and $Q$ heteroclinic connections $\gamma_{q}$. For any $\gamma_{q}$ and $x\in\gamma_{q}$ write $\alpha[\gamma_{q}]:=\alpha(x)$, $\omega[\gamma_{q}]:=\omega(x)$, so that $\gamma_{q}$ is a connection between the equilibria $\alpha[\gamma_{q}]$ and $\omega[\gamma_{q}]$. Itineraries along the network N can be encoded by sequences of transitions:
@@
Remark 2.2.
In some cases itineracy in terms of sequences of transitions $\mathbf{q}$ as in Definition 2.1 can be identified with sequences of equilibria: Suppose that for each pair $(\xi_{p},\xi_{p^{\prime}})$ there is at most one connecting trajectory $\gamma_{q}$ in N with $\alpha[\gamma_{q}]=\xi_{p}$ and $\alpha[\gamma_{q}]=\xi_{p^{\prime}}$. Then one may identify the sequence $\mathbf{q}=q_{1}q_{2}\ldots$ of connections with a sequence of equilibria $\mathbf{p}=p_{1}p_{2}\ldots\in\left\{1,\dotsc,P\right\}^{\mathbb{N}}$ of equilibria such that $\xi_{p_{k}}=\alpha[\gamma_{q_{k}}]$, $\xi_{p_{k+1}}=\omega[\gamma_{q_{k}}]$.
@@
an invariant set containing all connections in $\mathbf{q}$ as well as the equilibria that are their endpoints.
@@
A (homoclinic or heteroclinic) cycle as an invariant set C can be associated with a periodic sequence $\mathbf{q}$. In this case $\Gamma(\mathbf{q})=\textsf{{C}}$ is topologically a circle, because our definition of a cycle restricts attention to a single connection between a given pair of equilibria. Similarly, omnicycles are associated with periodic sequences. Conversely, for any periodic sequence $\mathbf{q}$ the invariant set $\Gamma(\mathbf{q})$ is, in general, topologically a union of circles.
Moving along a sequence now corresponds to a shift operation on a sequence of $Q$ symbols for each of the connections in the network N.
@@
The Kirk–Silber network [19] is the union of two distinct cycles with one joint connection between four equilibria $\xi_{1},\dotsc,\xi_{4}$; cf. Figure 1. If $\xi_{p}\to\xi_{p^{\prime}}]$ denotes the heteroclinic connection from $\xi_{p}$ to $\xi_{p^{\prime}}$, then the first cycle has connections $\gamma_{1}=[\xi_{1}\to\xi_{2}]$, $\gamma_{2}=[\xi_{2}\to\xi_{3}]$, $\gamma_{3}=[\xi_{3}\to\xi_{1}]$ and the second cycle has $\gamma_{1}=[\xi_{1}\to\xi_{2}],\gamma_{4}=[\xi_{2}\to\xi_{4}]$, $\gamma_{5}=[\xi_{4}\to\xi_{1}]$. The cycles correspond to the periodic sequences $123123\dotsb$ and $145145\dotsb$, respectively.
The only relevant information is what happens at the distribution node $\xi_{2}$: All sequences $\mathbf{q}$ that start with the joint connection $\gamma_{1}$ are concatenations of the symbols $\mathrm{A}:=123$, $\mathrm{B}:=145$, i.e., they can be written compactly as sequences on the symbols $\left\{\mathrm{A},\mathrm{B}\right\}$. (For sequences of equilibria $\mathbf{p}$, the symbols are $\mathrm{A}:=123$, $\mathrm{B}:=124$.) The shift $\Sigma$ induces a shift operator $\widetilde{\Sigma}:\left\{\mathrm{A},\mathrm{B}\right\}^{\mathbb{N}}\to\left\{\mathrm{A},\mathrm{B}\right\}^{\mathbb{N}}$ that corresponds to a full transition around one of the two cycles.
While there are only two heteroclinic cycles in this network, there are infinitely many omnicycles. These correspond to periodic sequences such as $\mathbf{q}=123145123145\dotsb$, which consists of cycle $\mathrm{A}$ once followed by $\mathrm{B}$ once.
$\xi_{1}$$\xi_{2}$$\xi_{3}$$\xi_{4}$$\gamma_{1}$$\gamma_{2}$$\gamma_{3}$$\gamma_{4}$$\gamma_{5}$$\mathrm{A}$$\mathrm{B}$
@@ 2.2. Switching near a network
Sequences may be used to encode the dynamics near a network $\textsf{{N}}\subset\mathbb{R}^{d}$. Let $U_{\textsf{{N}}}$ be a neighborhood of the network and let $U_{p}$ be a neighborhood of $\xi_{p}$. For each connection $\gamma_{q}$ in N consider a point $\zeta_{q}\in\gamma_{q}$ and let $V_{q}$ be a neighborhood of $\zeta_{q}$. Assume that all neighborhoods $U_{p},V_{q}$ are disjoint.
@@
[…]
@@
It follows from the definition above that $x\in\mathcal{S}_{\delta}(\mathbf{q})$ implies $x\in\mathcal{S}_{\delta}(\Sigma\mathbf{q})$. This can be seen by applying a suitable shift to the sequences of times $t_{k},t^{\prime}_{k}$. Hence, $\mathcal{S}_{\delta}(\mathbf{q})\subset\mathcal{S}_{\delta}(\Sigma\mathbf{q})$. In this sense, the sequence that is realized by a given trajectory near the network is unique only up to shifting. In particular, since $t_{k},t^{\prime}_{k}\geq 0$ above, a maximal sequence realized by a given trajectory can be obtained by checking near which connection this trajectory first enters $B_{\delta}(\textsf{{N}})$ permanently, i.e., it remains in $B_{\delta}(\textsf{{N}})$ for all future times.
Equivalenty, a single trajectory realizes a countable number of sequences on the network, all of which are related by the shift operator $\Sigma$. This number is finite if and only if the maximal sequence realized by this trajectory is preperiodic: By the argument above, $x\in\mathcal{S}_{\delta}(\mathbf{q})$ implies $x\in\mathcal{S}_{\delta}(\Sigma^{k}\mathbf{q})$ for all $k\in\mathbb{N}$. This means that if $x$ realizes the sequence $\mathbf{q}$, then $x$ realizes $\Sigma^{k}\mathbf{q}$ for all $k\in\mathbb{N}$. If $\mathbf{q}$ is aperiodic, then $\Sigma^{k}\mathbf{q}\neq\Sigma^{j}\mathbf{q}$ for $k\neq j$ and $x$ realizes the infinitely many distinct sequences $\Sigma^{k}\mathbf{q}$. If on the other hand $\mathbf{q}$ is preperiodic, then the set $\{\Sigma^{k}\mathbf{q}\mid k\in\mathbb{N}\}$ contains only finitely many distinct sequences, realized by $x$.
@@
There is (infinite) switching in positive time near the network N if in any $\delta$-neighborhood of N, every sequence $\mathbf{q}$ is realized by at least one trajectory, that is, for any $\delta>0$ and all $\mathbf{q}$ we have $\mathcal{S}_{\delta}({\mathbf{q}})\neq\emptyset$.
Remark 2.9.
Note that switching in positive time is sometimes also called forward switching. Analogously, one can consider backwards switching, or even bi-infinite switching. In this work we focus exclusively on forward switching and thus refer to it solely as switching.
@@
When restricting to equilibria with real eigenvalues, it was shown recently in [20] that for a large class of networks infinite switching is impossible. Even finite switching is restricted: All finite sequences of a given length $k$ are realized in a network only up to a maximal bound $k^{\mathrm{switch}}\in\mathbb{N}$, i.e., if all sequences of length $k$ are realized, then $k<k^{\mathrm{switch}}$.
Example 2.10.
We illustrate this restriction for the Kirk–Silber network (see Figure 1 and Example 2.5). Consider sequences of length three with $\gamma_{1}$ in the middle entry; the four possibilities that are compatible with the network architecture are $312$, $314$, $512$, $514$. In their original paper [19], Kirk and Silber already showed that for any given parameter configuration only three of these are realized. Thus, there is no switching near the Kirk–Silber network, not even along the common connection $\gamma_{1}$. In the framework of [20], this corresponds to $k^{\mathrm{switch}}=3$, that is there is a sequence of length three on the network that is not realized (which one it is depends on the parameters). By contrast, any sequence of length $k\leq 2$ is always realized, which implies switching at every node in the network.
@@ 3. Asymptotic stability of sequences
For a given sequence $\mathbf{q}$ on a network $\textsf{{N}}\subset\mathbb{R}^{d}$, we are interested in how many initial conditions $x$ lead to trajectories $\phi_{t}(x)$ that not only stay close to N but rather approach N as $t\to\infty$. From now on, let $\mu$ denote the Lebesgue measure on $\mathbb{R}^{d}$.
@@ 3.1. Fragmentary asymptotic stability of sequences
(3) $\mathcal{A}_{\delta}({\mathbf{q}})$ $:=\big\{x\in\mathcal{S}_{\delta}({\mathbf{q}})\,\big|\;\lim_{t\to\infty}\textnormal{d}(\textsf{{N}},\phi_{t}(x))=0\big\}.$
@@
Typically, stability is defined for sets which are (forward) invariant with respect to the flow $\phi_{t}$. Our set $\mathcal{A}_{\delta}({\mathbf{q}})$ is not forward invariant: The trajectory through a point $x\in\mathcal{A}_{\delta}({\mathbf{q}})$ follows $\mathbf{q}$ by definition. But the same is usually not true if we cut off a finite transient by considering $\phi_{t}(x)$ for $t>0$ sufficiently large, since $\phi_{t}(x)$ might be in an entirely different part of the network than the heteroclinic connection that is the first entry of $\mathbf{q}$. Thus, $\phi_{t}(x)\not\in\mathcal{A}_{\delta}(\mathbf{q})$, which means $\mathcal{A}_{\delta}(\mathbf{q})$ is not invariant. Applying the shift $\Sigma$, however, we can also remove the transient from $\mathbf{q}$, which means for some $\kappa(t)\in\mathbb{N}$ we have $\phi_{t}(x)\in\mathcal{A}_{\delta}(\Sigma^{\kappa(t)}\mathbf{q})$.
A concrete example for this can be seen in Figure 2 for $\mathbf{q}=134\dotsb$. Take a point $x$ on the blue trajectory in a $\delta$-neighborhood of $\alpha[\gamma_{1}]$, which is in $\mathcal{A}_{\delta}(134\dotsb)$. In forward time there is a $t>0$ such that $\phi_{t}(x)$ is in a $\delta$-neighborhood of $\omega[\gamma_{1}]=\alpha[\gamma_{3}]$. This means that
$\phi_{t}(x)\in\mathcal{A}_{\delta}(34\dotsb)=\mathcal{A}_{\delta}(\Sigma(134\dotsb)),$
but in general $\phi_{t}(x)\not\in\mathcal{A}_{\delta}(134\dotsb)$. In this case we have $\kappa(t)=1$.
This observation together with the desire for an invariant set that collects initial conditions with identical (asymptotic) behavior near the network prompts the next definition.
@@
(4) $\mathcal{B}_{\delta}({\mathbf{q}})$ $:=\left\{\,x\in B_{\delta}(\textsf{{N}})\,\left|\;\exists k\in\mathbb{N}\text{ such that }x\in\mathcal{A}_{\delta}(\Sigma^{k}\mathbf{q})\right.\right\}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{\delta}(\Sigma^{k}\mathbf{q}).$